OpenStudy (anonymous):

evaluate. log(base3) 9^500

OpenStudy (saifoo.khan):

\[\Huge \log_{3} 9^{500}\]

OpenStudy (saifoo.khan):

this?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

\[9^{500}=3^{1000}\] so answer is 1000

OpenStudy (anonymous):

only step is writing \[9^{500}=(3^2)^{500}=3^{1000}\]

OpenStudy (anonymous):

and \[\log_3(3^x)=x\]

OpenStudy (anonymous):

easy thanks guys